OpenStudy (anonymous):

on the 4 major or math tests given this term, awilda had grades of 78,84,69,83 she would like her test average to 80 what mark must awilda get on her next test

OpenStudy (zzr0ck3r):

\[\frac{78+84+69+83+t}{5}=80\]

OpenStudy (zzr0ck3r):

solve for t

OpenStudy (anonymous):

ok

OpenStudy (zzr0ck3r):

do you understand who I got that equation?

OpenStudy (anonymous):

yea i guess

OpenStudy (zzr0ck3r):

you guess?

OpenStudy (anonymous):

do u divide?

OpenStudy (zzr0ck3r):

if we average 5 numbers then \[\frac{a+b+c+d+e}{5}=average\] in your case we are given the first 4 numbers, we need to find the 5th

OpenStudy (anonymous):

so add the top and divide it by 5?

OpenStudy (zzr0ck3r):

\[\frac{78+84+69+83+t}{5}=80\] multiply both sides by 5 \[78+84+69+83+t = 400\]

OpenStudy (zzr0ck3r):

can you solve the rest?

OpenStudy (anonymous):

ok im going to solve it

OpenStudy (anonymous):

80?

OpenStudy (zzr0ck3r):

314+t = 400

OpenStudy (zzr0ck3r):

now can you solve it?

OpenStudy (anonymous):

86 i meant

OpenStudy (littlebird):

You get an average by adding the numbers together and dividing by the number of integers. In this case, there are 5 numbers. You have to add them together and divide them by 5 with the end result being 80.

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

thanks man!

OpenStudy (zzr0ck3r):

np